# Dissociation of Faithful and Unfaithful Reasoning in LLMs

### v1

@@ Abstract
Large language models (LLMs) improve their performance in downstream tasks when they generate Chain of Thought reasoning text before producing an answer. Our research investigates how LLMs recover from errors in Chain of Thought, reaching the correct final answer despite mistakes in the reasoning text. Through analysis of these error recovery behaviors, we find evidence for unfaithfulness in Chain of Thought, but we also identify many clear examples of faithful error recovery behaviors. We identify factors that shift LLM recovery behavior: LLMs recover more frequently from obvious errors and in contexts that provide more evidence for the correct answer. However, unfaithful recoveries show the opposite behavior, occurring more frequently for more difficult error positions. Our results indicate that there are distinct mechanisms driving faithful and unfaithful error recoveries. Our results challenge the view that LLM reasoning is a uniform, coherent process.
Evelyn Yee,   Alice Li  [^1] Equal contribution,   Chenyu Tang*,   Yeon Ho Jung [^2] Work done at University of California, San Diego.
Ramamohan Paturi [^3] Equal contribution,   Leon Bergen  [^4] Correspondence to lbergen@ucsd.edu
University of California, San Diego
@@ 1 Introduction
Large language models have shown a remarkable ability to solve high-level reasoning tasks across different domains (Bubeck et al., 2023; Burnell et al., 2023; Chang & Bergen, 2023; Yu et al., 2023). Chain of thought (CoT), a prompting strategy that involves breaking down complex tasks into smaller sub-tasks and using the outputs of one sub-task as the input for the next, has been crucial for improvements in model performance on these reasoning tasks (Wei et al., 2023). In both the few-shot and zero-shot prompting styles, chain of thought improves language model performance on multi-step reasoning, problem-solving, and Q&A tasks (Kojima et al., 2022; Suzgun et al., 2022; Wei et al., 2023) in English and multilingual contexts (Shi et al., 2022). Additionally, chain of thought has been shown to work in multimodal (text+image) contexts (Rose et al., 2023; Zhang et al., 2023). More generally, there may be fundamental reasons which explain the importance of chain of thought for language model reasoning (Feng et al., 2023).
Despite the significant empirical success of chain of thought and the beginnings of a theoretical understanding, there is still much unknown about the mechanistic reasons for its success (Madaan & Yazdanbakhsh, 2022; Wang et al., 2023). Although chain of thought transcripts resemble human reasoning on a surface level (Dziri et al., 2023; Wei et al., 2023), it is unknown whether this observed output aligns faithfully with the language model’s true internal reasoning processes. This question of ”faithfulness” (Jacovi & Goldberg, 2020) in chain of thought is fundamental for understanding how language models reason, and whether their reasoning is a trustworthy source of information for human users.
In this study, we investigate the faithfulness of chain of thought by focusing on language models’ ability to recover from errors in their chain of thought texts. Our study makes use of the dissociation paradigm from psychology and neuroscience (Tulving, 1972; Shallice, 1988). We apply interventions to LLM reasoning and measure the effect on faithful and unfaithful error recoveries. If these behaviors emerge differently, this provides evidence for distinct mechanisms for faithful and unfaithful error recoveries.
@@ 2.1 Understanding chain of thought
Language models have been observed to naturally produce errors in chain of thought generation, and there have been some efforts to describe these errors. For example, Zhang et al. (2023) categorized chain of thought errors in a multi-modal system as ”hallucination” or ”other.” Similarly, Nguyen et al. (2024) classify chain of though errors into factual errors, coherence errors, and answer errors, noting a discrepancy between chain of thought accuracy and answer accuracy, which may grow with model size. McKenna et al. (2023) presented an investigation of reasoning errors in LLMs and attributed hallucination errors to memorization from their training corpus. Dziri et al. (2023) provided more fine-grained categorizations of errors in generated reasoning chains, identifying that local errors can either propagate or be resolved in later reasoning steps. That work also provided theoretical arguments that the propagation effect should significantly overwhelm the recovery effect, making errors inevitable for arbitrarily long reasoning chains. This effect was similarly observed by Arora et al. (2022), which developed a theoretical model for error propagation in model-generated text.
@@
Recent research points to instances of unfaithful behavior in chain of thought. Agarwal et al. (2024) distinguish ”plausible” explanations from ”faithful” ones and call for further research and development of both plausible and faithful systems. Turpin et al. (2023) showed that language models can use chain of thought to unfaithfully rationalize answers that are derived from superficial cues in the prompt. To a similar effect, Lanham et al. (2023) implemented counterfactual interventions on chain of thought texts, finding that LLMs can achieve the correct answer despite variations on their chain of thought, like early stopping and error introductions. Gao (2023) also used counterfactual interventions, similar to those in the current study, and employed Shapely values to assess the significance of tokens in the model’s reasoning text. Importantly, both studies found that, despite logically disrupted reasoning, the model’s conclusions often remained unaffected. This suggests the model’s final answer is not always aligned with its generated reasoning text. To further understand this effect, Bao et al. (2024) identify the ”instructions” of a task as a potential mediating factor in the disconnect between the chain of thought text and the model’s final answer.
@@ 3 Methods
In all of our experiments, we measure the effect of introducing errors into chain of thought reasoning text. These errors are generated in several stages (see Figure 1). First, given a question from a dataset, an LLM is prompted with Kojima et al. (2022)’s zero-shot CoT prompting method (“Let’s think step-by-step”). If the answer generated at the end of the CoT text is correct, this question and CoT text is kept; if the answer is incorrect, the question and CoT text is removed. The CoT texts that resulted in correct answers make up the ground-truth CoT transcripts.
Given a ground-truth CoT transcript, a number in the text is selected, and an error is introduced at this point. The LLM is provided the question and the CoT text up to the point of this error and is asked to complete the reasoning text. In general, selected values were adjusted by random integer values in {-3, -2, -1, 1, 2, 3}, following the methodology of Gao (2023). Figure 2 shows an example transcript, and Appendix A.3 contains more examples for each error position.
@@ 3.1 Models
We evaluated fixed versions of GPT 3.5 (gpt-3.5-turbo-0301) and GPT-4 (gpt-4-0314) OpenAI (2023). For the primary analysis, we focus on GPT-4 results, but GPT-3.5 results are available in the Appendix. All model responses were gathered using the OpenAI public API, and sampling was performed using greedy decoding (temperature = 0).
@@ 3.2 Datasets
For each model in each dataset, we randomly sampled 300 $<$question, chain of thought, answer$>$ triples for which the model achieved the correct answer, indicating that the original chain of thought was correct. [^1] A sample size of 300 was selected due to the costs of manual annotation. These triples, which were collected separately for each model and dataset, make up our ground-truth data for all further experiments.
3.3 Stimulus Validation
Numerical errors were introduced using regular expressions. For all experiments, we manually verified that the errors introduced were of the correct type. We also verified that the errors were essential to the logic of the problem solution, to guarantee that correct final answers truly indicate error recovery (as opposed to the error having no effect on the final answer if naively propagated through). This resulted in slightly varying sample sizes across experimental conditions.
3.4 Faithfulness Annotation
For more details on our annotation process, including stimulus validation and faithfulness annotation, see Appendix A.4.
3.5 Statistical Analyses
We use multinomial logistic regression with fixed effects for datasets in order to estimate effects of different variables on error recovery. [^2] Random effects models did not converge in this setting.
@@ 4 Error recovery behaviors
Several previous studies have observed error recovery as a general phenomenon in chain of thought. (See Section 2 for an overview.) Generally, it is assumed that error recovery indicates unfaithfulness, as the model reaches an error-free answer despite errors in the reasoning text. However, it is not clear whether recovery always indicates unfaithfulness.
@@
5 Experiment 1: Error evidence
@@
Experiment 1 is motivated by the following observation: different types of errors will provide the model with different amounts of evidence regarding how to correctly recover. If the error is introduced into information that the model has previously seen or computed, then it can identify that an error has occurred by noticing a discrepancy with the prior information, and recover by copying the old information into its current context. In contrast, if the error is propagated through the CoT text, occurring multiple times, then the model will have stronger evidence that no error has occurred, and that its prior reasoning is in fact correct.
@@
We hypothesize that error recovery will occur more often when an LLM has more evidence about how to correctly recover. We investigate this hypothesis by introducing errors on three distinct positions in the CoT text. See Figure 3 for an illustration.
@@ 5.1 Results
See Appendix B.1 for full results from this experiment, including evaluation of GPT-3.5.
Figure 4: Total error recovery rates for GPT-4 for each error position. Error bars indicate 95% binomial confidence intervals.
Figure 4 shows the overall error recovery results for GPT-4 for the different error positions. GPT-4 had the highest error recovery rate in the copying error condition and the lowest error recovery rate in the propagated calculation error condition ($p<0.001$ for both comparisons, from multivariate logistic regression).
5.1.1 Faithful and unfaithful recoveries
Figure 5: Faithful and unfaithful recovery rates for GPT-4. Error bars indicate 95% binomial confidence intervals.
@@ 5.1.1 Faithful and unfaithful recoveries
A different pattern was observed for unfaithful recoveries. Compared to faithful recoveries, there was a significantly smaller effect of calculation errors ($p<0.005$) or propagated calculation errors ($p<0.001$) on the rate of unfaithful recoveries.
@@
6 Experiment 2: Error magnitude
In Experiment 2, we manipulate the perceptibility of errors by changing their magnitude. Errors with greater magnitude are expected to be more noticeable to the model, resulting in higher rates of recovery.
The ground truth chain of thought datasets and procedure for introducing errors are the same as Experiment 1. However, rather than perturbing the selected values by a random amount, as in Experiment 1, our perturbations are more controlled.
6.1 Results
See Appendix B.2 for full results from this experiment. Overall, error recovery occurred more often for large magnitude errors than small ones ($p<0.001$), as shown in Figure 11.
Figure 6: Difference between large error and small error recovery rates for GPT-4. Negative values indicate recoveries occurred more often for small errors. Error bars are 95% confidence intervals.
Figure 6 shows the effect of error magnitude on faithful and unfaithful recoveries. Large errors led to higher rates of faithful recovery than small errors ($p<0.001$). However, the effect is reversed for unfaithful recovery. The rate of unfaithful recovery was higher for small errors than large errors, and there is a significant difference between the effects for faithful and unfaithful recovery ($p<0.001$).
The results in this experiment provide evidence that GPT-4 faithfully recovers more frequently from large errors than small errors, suggesting that faithful recovery becomes easier when errors are more noticeable. However, the opposite was found for unfaithful recoveries. These recoveries occurred more often when errors were smaller and harder to notice.
7 Experiment 3: Prior expectations
Experiment 3 evaluates the hypothesis that a model will recover more frequently if it expects that an error is likely to occur in its CoT transcript. We increase this prior expectation of an error using two methods: introducing noise into the transcript, or directly prompting the model with this information.
Error Recovery Prompt To more explicitly induce error expectations in the model, we modified the chain of thought prompt. In all of other evaluations, we prefaced the CoT transcript with the sentence, ”Let’s think step by step.” (Kojima et al., 2022). In contrast, in this condition, we remind the model to specifically look for errors, replacing the previous prompt with ”Let’s think step by step, being careful to notice and fix any mistakes.”
7.1 Results
Figure 14 shows overall recovery behavior. Context noise increased the recovery rate relative to baseline ($p<0.001$), and the error recovery prompt had a smaller but still significant effect ($p<0.001$).
Figure 7: Difference in GPT-4’s recovery rates between context noise and baseline conditions. Negative values indicate recoveries occurred more often in the baseline condition. Error bars are 95% confidence intervals.
Figure 7 shows the effect of context noise on faithful and unfaithful recoveries. Introducing context noise increased the rate of faithful recovery and decreased the rate of unfaithful recovery. The difference between faithful and unfaithful recovery was significant for context noise ($p<0.001$). The error recovery prompt had a smaller effect on both faithful and unfaithful recovery rates (Figure 17), so no difference between these effects could be detected ($p=0.42$).
See Appendix B.3 for full results from this experiment.
The results provide evidence that GPT-4 recovers from errors more frequently when the context provides evidence that errors will occur in the CoT transcript. The effect is specific to faithful recoveries, with unfaithful recoveries showing the opposite pattern.

### v2

@@ Abstract
Large language models (LLMs) often improve their performance in downstream tasks when they generate Chain of Thought reasoning text before producing an answer. We investigate how LLMs recover from errors in Chain of Thought. Through analysis of error recovery behaviors, we find evidence for unfaithfulness in Chain of Thought, which occurs when models arrive at the correct answer despite invalid reasoning text. We identify factors that shift LLM recovery behavior: LLMs recover more frequently from obvious errors and in contexts that provide more evidence for the correct answer. Critically, these factors have divergent effects on faithful and unfaithful recoveries. Our results indicate that there are distinct mechanisms driving faithful and unfaithful error recoveries. Selective targeting of these mechanisms may be able to drive down the rate of unfaithful reasoning and improve model interpretability.
@@ 1 Introduction
Large language models have shown a remarkable ability to solve high-level reasoning tasks across different domains (Bubeck et al., 2023; Burnell et al., 2023; Chang & Bergen, 2023; Yu et al., 2023). Chain of thought (CoT), a prompting strategy that involves breaking down complex tasks into smaller sub-tasks and using the outputs of one sub-task as the input for the next, has been crucial for improvements in model performance on these reasoning tasks (Wei et al., 2023). In both the few-shot and zero-shot prompting styles, chain of thought improves language model performance on multi-step reasoning, problem-solving, and Q&A tasks (Kojima et al., 2022; Suzgun et al., 2022; Wei et al., 2023) in English and multilingual contexts (Shi et al., 2022). Additionally, chain of thought has been shown to work in multimodal contexts (Rose et al., 2023; Zhang et al., 2023). More generally, there may be fundamental reasons which explain the importance of chain of thought for language model reasoning (Feng et al., 2023).
Despite the significant empirical success of chain of thought and the beginnings of a theoretical understanding, there is still much unknown about the mechanistic reasons for its success (Madaan & Yazdanbakhsh, 2022; Wang et al., 2023). Although chain of thought transcripts resemble human reasoning on a surface level (Dziri et al., 2023; Wei et al., 2023), it is unknown whether this observed output aligns with the language model’s internal reasoning processes.
The current study investigates the conditions under which LLM reasoning is faithful, where reasoning text is considered faithful if it provides a valid or reasonable argument in support of the final conclusion. This question of ”faithfulness” (Jacovi & Goldberg, 2020) in chain of thought is fundamental for understanding whether their reasoning is a trustworthy source of information for human users. Faithful reasoning ensures that the model’s output not only reaches the correct conclusion but does so in a logically valid manner, allowing for human verification.
We investigate the faithfulness of chain of thought by focusing on language models’ ability to recover from errors in their chain of thought texts. The study makes use of the dissociation paradigm from psychology and neuroscience (Tulving, 1972; Shallice, 1988). We apply interventions to LLM reasoning and measure the effect on faithful and unfaithful error recoveries. If these behaviors respond differently to the interventions, this provides evidence for distinct mechanisms for faithful and unfaithful error recoveries.
@@ 2.1 Understanding chain of thought
Language models have been observed to produce errors in chain of thought generation, and there have been some efforts to describe these errors. For example, Zhang et al. (2023) categorized chain of thought errors in a multi-modal system as ”hallucination” or ”other.” Similarly, Nguyen et al. (2024) classify chain of though errors into factual errors, coherence errors, and answer errors, noting a discrepancy between chain of thought accuracy and answer accuracy, which may grow with model size. McKenna et al. (2023) presented an investigation of reasoning errors in LLMs and attributed hallucination errors to memorization from their training corpus. Dziri et al. (2023) provided more fine-grained categorizations of errors in generated reasoning chains, identifying that local errors can either propagate or be resolved in later reasoning steps. That work also provided theoretical arguments that the propagation effect should significantly overwhelm the recovery effect, making errors inevitable for arbitrarily long reasoning chains. This effect was similarly observed by Arora et al. (2022), which developed a theoretical model for error propagation in model-generated text.
@@
Recent research points to instances of unfaithful behavior in chain of thought. Agarwal et al. (2024) distinguish ”plausible” explanations from ”faithful” ones and call for further research and development of both plausible and faithful systems. Turpin et al. (2023) showed that language models can use chain of thought to unfaithfully rationalize answers that are derived from superficial cues in the prompt. To a similar effect, Lanham et al. (2023) implemented counterfactual interventions on chain of thought texts, finding that LLMs can achieve the correct answer despite variations on their chain of thought, like early stopping and error introductions. Gao (2023) also used counterfactual interventions, similar to those in the current study, and employed Shapley values to assess the significance of tokens in the model’s reasoning text. Importantly, both studies found that, despite logically disrupted reasoning, the model’s conclusions often remained unaffected. This suggests the model’s final answer is not always aligned with its generated reasoning text. To further understand this effect, Bao et al. (2024) identify the ”instructions” of a task as a potential mediating factor in the disconnect between the chain of thought text and the model’s final answer.
Additional works examine faithfulness in similar textual reasoning contexts, including Free-text Rationales (Wiegreffe et al., 2021) and Natural Language Explanations (Atanasova et al., 2023). Across both CoT and NLE contexts, Parcalabescu & Frank (2024) remark that existing faithfulness tests can be framed as measuring output consistency rather than internal model functioning.
@@ 3 Methods
In our experiments, we measure the effect of introducing errors into chain of thought reasoning text. These errors are generated in several stages (see Figure 1). First, given a question from a dataset, an LLM is prompted with Kojima et al. (2022)’s zero-shot CoT prompting method (“Let’s think step-by-step”). If the CoT text is logically valid (i.e. does not contain unnecessary steps, mistakes, or inconsistent reasoning) and the answer generated at the end of the CoT text is correct, this question and CoT text are kept; if the reasoning is invalid or the answer is incorrect, the question and CoT text are removed. This filtering is done manually. The resulting valid, correct CoT texts make up the ground-truth CoT transcripts.
Given a ground-truth CoT transcript, a number in the text is selected, and an error is introduced at this point. The LLM is provided with the question along with the previous CoT text up to the point of this error, and it responds by completing the reasoning text. In general, selected values were adjusted by random integer values in {-3, -2, -1, 1, 2, 3}, following the methodology of Gao (2023). Figure 2 shows an example transcript.
@@ 3.1 Models
We tested fixed versions of OpenAI’s GPT-4 OpenAI (2023) as well as Anthropic’s Claude-3 Opus and Meta’s Llama-3 70B Chat model. The GPT model responses were gathered using the OpenAI API (gpt-3.5-turbo-0301 and gpt-4-0314), Claude responses were gathered using the Anthropic API (claude-3-opus-20240229), and LLama responses were gathered using the Together AI API (meta-llama/Llama-3-70b-chat-hf). Decoding was done with temperature 0.
@@ 3.2 Datasets
For each model in each dataset, we randomly sampled 300 $<$question, chain of thought, answer$>$ triples for which the model achieved the correct answer. [^1] A sample size of 300 per dataset was selected based on a statistical power analysis, to maximize sensitivity while reducing the costs of manual annotation. These triples, which were collected separately for each model and dataset, make up our ground-truth data for all further experiments.
@@ 4 Error recovery behaviors
4.1 Error Introduction
Numerical errors were introduced using regular expressions. In the context of math word problems, we targeted errors in critical calculation steps, which derive a new numerical value which is essential to achieving the correct final answer. For all experiments, we manually verified that the errors met these criteria, to guarantee that correct final answers truly indicated error recovery, as opposed to the error having no effect on the final answer if naively propagated through. This resulted in slightly varying sample sizes across experimental conditions.
4.2 Faithfulness Analysis
For more details on our annotation process, including stimulus validation and faithfulness annotation criteria, see Appendix A.4.
In order to estimate the effects of different experimental variables on error recovery and faithfulness, we use multinomial logistic regression with fixed effects for datasets. [^2] Random effects models did not converge in this setting.
Several previous studies have observed error recovery, where a model reaches a correct final answer despite a flaw in the intermediate reasoning, as a general phenomenon in chain of thought. (See Section 2 for an overview.) Generally, it is assumed that error recovery indicates unfaithfulness, as the model reaches an error-free answer despite errors in the reasoning text. However, it is not clear whether recovery always indicates unfaithfulness. For example, it could be possible that a model explicitly announces the presence of an error and state a plausible, complete process for recovery. In this case, the error recovery would present no evidence for unfaithfulness.
@@ 4.2 Faithfulness Analysis
5 Experiment 1: Error magnitude
Figure 3: Overall error recovery rates (as a proportion of all responses) from small errors and large errors. Error bars indicate 95% binomial confidence intervals.
In Experiment 1, we manipulate the perceptibility of errors by changing their magnitude (i.e. the absolute numerical difference between the error value and the original value). Larger errors are expected to be more noticeable to the model, resulting in higher rates of recovery.
5.1 Results
Figure 4: Difference between large error and small error recovery rates, as a proportion of all responses. Negative values indicate recoveries occurred more often for small errors. Error bars are 95% confidence intervals.
See Appendix C.1 for full results from this experiment.
As shown in 3, the different models’ error recovery capabilities varied. All models showed higher error recovery rates for large magnitude errors. ($p<0.001$ for each model).
Figure 4 shows the effect of error magnitude on faithful and unfaithful recoveries. For GPT-4, increased error size led to higher rates of faithful recovery and lower rates of unfaithful recovery. For Claude-3 Opus, it led to a small decrease in faithful error recovery and a large increase in unfaithful recovery. For Llama-3, it led to a small increase in faithful error recovery and a large increase in unfaithful error recovery. For all three models, we found the faithful response to error magnitude and the unfaithful response to error magnitude to be significantly different ($p<0.001$ for GPT-4 and Llama-3, and $p<0.05$ for Claude-3 Opus).
The results in this experiment provide evidence that large language models recover more frequently from large errors than small ones. Consistent across models of different families, we find a dissociation between faithful and unfaithful modes of error recovery, though the individual responses vary between models.
6 Experiment 2: Prior expectations
Experiment 2 evaluates the hypothesis that a model will recover more frequently if it expects that an error is likely to occur in its CoT transcript. We increase this prior expectation of an error using two methods: introducing noise into the transcript, or directly prompting the model with this information.
Error Recovery Prompt To more explicitly induce error expectations in the model, we modified the chain of thought prompt. In all of the other experiments, we prefaced the CoT transcript with the sentence, ”Let’s think step by step.” (Kojima et al., 2022). In contrast, in this condition, we remind the model to specifically look for errors, replacing the previous prompt with ”Let’s think step by step, being careful to notice and fix any mistakes.”
Figure 5: Difference in recovery rates (as a proportion of all responses) between context noise and baseline conditions. Negative values indicate recoveries occurred more often in the baseline condition. Error bars are 95% confidence intervals.
6.1 Results
See Appendix C.2 for full results from this experiment.
Figure 12 shows overall recovery behaviors. Context noise increased the recovery rate relative to baseline ($p<0.001$ for all models). The error recovery prompt had a less pronounced effect, producing a small increase in total error recovery from GPT-4 ($p<0.05$) and Claude-3 Opus ($p<0.001$) and no detectable change in total error recovery from Llama-3 ($p>0.05$). Our analyses focus primarily on the effect of context noise.
Figure 5 shows the effect of context noise on faithful and unfaithful recoveries. For GPT-4, introducing context noise increased the rate of faithful recovery and decreased the rate of unfaithful recovery. For Claude-3 Opus and Llama-3, we observed that context noise led to an increase in unfaithful recovery and a slight decrease in faithful recovery. We found this difference between faithful and unfaithful recovery to be significant for GPT-4 ($p<0.001$) and Claude-3 Opus ($p<0.001$), but not for Llama-3 ($p=0.07$), due to the smaller overall response to context noise in Llama-3.
As shown in Figure 15, the error recovery prompt led to increased faithful recovery and generally decreased unfaithful recovery for all 3 models. This difference between faithful and unfaithful behavior was not significant for GPT-4 ($p=0.08$), but it was significant for Claude-3 Opus ($p<0.001$) and Llama-3 ($p<0.001$).
The results provide evidence that language models recover from errors more frequently when the context provides evidence that errors will occur in the CoT transcript. Across all models, this contextual evidence has distinct effects on faithful recoveries compared to unfaithful recoveries. However, the direction of the effects varied across the contextual manipulations and models.
7 Experiment 3: Recoverability
@@
Experiment 3 is motivated by the following observation: different types of errors will provide the model with different amounts of evidence regarding how to correctly recover. If the error is introduced into information that the model has previously observed, then it can identify that an error has occurred by noticing a discrepancy with the prior information, and recover by copying the old information into its current context. In contrast, if the error is propagated through the CoT text, occurring multiple times, then the model will have stronger evidence that no error has occurred and that its prior reasoning is correct.
@@
We hypothesize that error recovery will occur more often when an LLM has more evidence about how to correctly recover. We investigate this hypothesis by introducing errors on three distinct positions in the CoT text. See Figure 6 for an illustration Appendix A.3 for more examples for each error position.
@@ 7.1 Results
Figure 16 shows the overall error recovery results for GPT-4, which was the only model evaluated in this experiment. See Appendix C.3 for full results from this experiment. GPT-4 had the highest error recovery rate in the copying error condition and the lowest error recovery rate in the propagated calculation error condition ($p<0.001$ for both comparisons).
Figure 7: Faithful and unfaithful recovery rates (as a proportion of all responses) for GPT-4 for each error position. Error bars indicate 95% binomial confidence intervals.
@@
A different pattern was observed for unfaithful recoveries. Compared to faithful recoveries, there was a significantly smaller effect of calculation errors ($p<0.05$) or propagated calculation errors ($p<0.001$) on the rate of unfaithful recoveries.